OpenStudy (anonymous):

f(x)=[(x+3)/2][x^2-4x+9] I need to find its derivative using product rule , Please Help !

OpenStudy (mrnood):

the product rule d/dx u.v = u dv/dx +v du/dx in your case: u= (x+3)/2 so find du/dx v= (x^2-4x+9) so find dv/dx then put all those bits into the equation above and simplify..

OpenStudy (anonymous):

du/dx = 1/2 and dv/dx = 2x-4 is this right?

OpenStudy (mrnood):

yes